# Extreme resonant eccentricity excitation of stars around merging black-hole binary

### v1

@@ Abstract
We study the dynamics of a star orbiting a merging black-hole binary (BHB) in a coplanar triple configuration. During the BHB’s orbital decay, the system can be driven across the apsidal precession resonance, where the apsidal precession rate of the stellar orbit matches that of the inner BHB. As a result, the system gets captured into a state of resonance advection until the merger of the BHB, leading to an extreme eccentricity growth of the stellar orbit. This resonance advection occurs when the inner binary has a non-zero eccentricity and unequal masses. The resonant driving of the stellar eccentricity can significantly alter the hardening rate of the inner BHB, and produce observational signatures to uncover the presence of nearby merging or merged BHBs.
@@
BHB progenitors, potentially numerous yet undetected, could form an abundant population in stellar multiples (triples or quadruples) Tokovinin ; Raghavan ; Fuhrmann ; Moe-2017 . Observationally, triple systems with massive stars are suggested to be less inclined or nearly coplanar, especially if the tertiary is not so faraway Ransom ; Thompson ; Eisner . If the tertiary companion is a visible object (e.g., a star or a pulsar), its motion can aid in the search for inspiral BHBs Suto-1 ; Suto-2 ; Suto-3 ; Liu-Daniel . In this letter, we focus on the secular dynamics of a star orbiting a merging BHB in a coplanar setup. Beginning with an eccentric BHB and a nearly circular outer stellar orbit, we demonstrate that the system can enter an apsidal resonance and be captured into a ‘resonance advection’ state during the inner BHB’s gravitational radiation driven orbital decay. Once captured, the eccentricity of the star grows continually, approaching unity in magnitude.
@@
Figure 1: The evolution of a merging BHB with a test-mass tertiary in a coplanar orbital configuration. The masses of BHB are $m_{1}=46.15M_{\odot}$ and $m_{2}=13.85M_{\odot}$. The initial semimajor axes and eccentricities of the inner and outer binaries are $a_{\mathrm{in},0}=0.4\mathrm{AU}$, $e_{\mathrm{in},0}=0.8$, $a_{\mathrm{out}}=5\mathrm{AU}$ and $e_{\mathrm{out},0}=0.1$, and the longitudes of the periapsis of the two orbits are set to be $\varpi_{\mathrm{in},0}=\varpi_{\mathrm{out},0}=0$. The black (inner binary) and orange (outer binary) lines represent the numerical solutions of the DA secular equations. Key moments during the inner BHB evolution are highlighted in the top two panels (gray dots) with the orbital parameter values labeled. The bottom panel displays the apsidal precession rates given by Eqs. (2)-(3).
@@
The top two panels of Fig. 1 show an example of the orbital evolution of a coplanar triple. The system satisfies the dynamical stability criterion Holman . The results are obtained by integrating the double-averaged (DA; averaging over both the inner and outer orbital periods) secular equations of motion Liuetal-2015 . All Newtonian effects (up to the octupole order), the leading order PN corrections, and the dissipation due to gravitational wave (GW) are included in the calculation. We see that the BHB undergoes orbital decay and eccentricity circularization due to GW radiation. The merger time of the BHB is approximately given by $T_{\mathrm{merger}}=[5c^{5}a_{\mathrm{in},0}^{4}(1-e_{\mathrm{in},0}^{2})^{7/2}]/[256G^{3}m_{12}^{2}\mu_{\mathrm{in}}]$ Peters-1964 .
@@
and $\dot{\varpi}_{\mathrm{out}}$ is due to the secular perturbation of the inner binary on the outer orbit Liuetal-2015 ; Liu-SMBHB
@@
We see that starting from a quasi-circular orbit, the outer eccentricity can be excited once $\dot{\varpi}_{\mathrm{out}}$ catches up to $\dot{\varpi}_{\mathrm{in}}$. Importantly, rather than a single resonance passage, both frequencies evolve in lockstep until the inner BHB merges, resulting in a significant growth in $e_{\mathrm{out}}$. It’s worth noting that the final $e_{\mathrm{out}}$ becomes quite large, and the DA approximation may be challenged. We employ alternative single-averaged equations of motion to assess its validity in Appendix A.
Figure 2: Phase-space portraits in terms of the canonical variables ($1-\sqrt{1-e_{\mathrm{out}}^{2}}$ and $\Delta\varpi$) at different stages of the orbital evolution for the system in our fiducial model (refer to Fig. 1). In each panel, $\gamma\equiv\dot{\varpi}_{\mathrm{in}}/\dot{\varpi}_{\mathrm{out}}|_{e_{\mathrm{out}}=0}$ and $\dot{\varpi}_{\mathrm{in}}/\dot{\varpi}_{\mathrm{out}}$ are evaluated by Eqs. (13) and (2)-(3) at the time labeled in Fig. 1. The solid gray lines are the contours of constant $\hat{\mathcal{H}}$ (see Eq. Extreme resonant eccentricity excitation of stars around merging black-hole binary). The existence of a separatrix (black line) divides phase-space into circulation zones (zone I and III) and libration zone (zone II). The orange curve represents the numerical solution of the outer binary’s eccentricity vector $\bm{e}_{\mathrm{out}}$, originating at the red dot and ending at the blue dot. To better visualize both evolution trajectory and separatrix, zoom-in subfigures are included in the top-left panels.
@@
By introducing $\Phi_{0}=\Phi|_{a_{\mathrm{in}}=a_{\mathrm{in},0}}$ with $a_{\mathrm{in},0}$ being the initial semimajor axis, Eq. (8) can be made dimensionless as
@@
Note that apsidal precession resonance and the possibility of exciting eccentricity at resonance haven been well recognized in the planetary dynamics literatures, often in the linear eccentricity regime (e.g., Heppenheimer ; Ward ; Lin ; XuWenrui ; Diego ; Suyubo ). However, the distinct advection effect, sweeping through the full range of canonical momentum, is a unique feature of our present study. Three requirements for the system to end up captured into resonance advection have to be satisfied Quillen ; Henrard , including:
@@
iii) Sufficiently small initial eccentricity. The initial eccentricity of the tertiary must be small (e.g., $e_{\mathrm{out},0}\lesssim 0.6$ in our case), so that it can be captured into the resonance while zone III has not yet appeared (see below and Appendix C).
Fig. 3 illustrates the excitation of stellar eccentricity for a range of $(e_{\mathrm{in},0},e_{\mathrm{out},0})$. We find that the cases with $e_{\mathrm{in},0}\gtrsim e_{\mathrm{out},0}$ always give rise to resonant driving of $e_{\mathrm{out}}$, with weak dependence on the initial eccentricities. This behavior can be understood as follows. In the absence of separatrix crossing, as noted before, the phase-space area $\mathcal{A}$ (Eq.14) is an adiabatic invariant. At the time of separatrix crossing, $\mathcal{A}$ is not conserved, and the trajectory can transition into nearby zones Suyubo ; Henrard . For a system initially set with a quasi-circular outer orbit, the trajectory of $\bm{e}_{\mathrm{out}}$ circulates near the bottom part of the phase-space portrait (see the top left panel of Fig. 2). Because there are only two zones at separatrix crossing, the system is guaranteed to be captured into zone II (as in the top panels of Fig. 2). When $0.3\lesssim e_{\mathrm{out}}\lesssim 0.6$, zone III emerges in the transition, and both zone II and III become available. The probabilities of transition to zone $\rm II$ and $\rm III$ are given by
@@
where $\mathcal{A}_{\rm I(\rm II,III)}$ is the area of the individual zone. Note that $d\mathcal{A}_{\rm I}+d\mathcal{A}_{\rm II}+d\mathcal{A}_{\rm III}=0$. In our case, both $P_{\rm I\rightarrow\rm II}$ and $P_{\rm I\rightarrow\rm III}$ are positive due to the expansion of zone II and zone III and the contraction of zone I at the separatrix crossing. Whether the system transitions to zone II or zone III depends on the initial $\Delta\varpi_{0}$, which affects the initial phase-space area for a given $e_{\mathrm{out},0}$ (see more details in the Appendix B). If $e_{\mathrm{out},0}$ is even higher ($\gtrsim 0.6$), with zone II starting to shrink, capture into zone II becomes impossible as $P_{\rm I\rightarrow\rm II}$ is negative.
Figure 3: The evolution of the outer binary eccentricity as a function of the inner binary semimajor axis. We consider different combinations of ($e_{\mathrm{in},0}$, $e_{\mathrm{out},0}$) in our fiducial model (with the parameters the same as in Fig. 1). All systems satisfy the stability criterion.
@@
Finite $m_{3}$: modified evolution of merging BHBs.— We now consider the ‘BHB$+$star’ system where the tertiary companion has a finite mass $m_{3}=1M_{\odot}$. In this case, the perturbation from the tertiary companion on the inner binary cannot be ignored. Since the outer binary is allowed to ‘gain’ some eccentricity from the inner BHB at resonance, the inner binary can experience an eccentricity decrease, therefore reducing the orbital decay rate of the BHB. When the initial $e_{\mathrm{out}}$ is large, the inner binary can ‘gain’ some eccentricity, therefore increasing the decay rate.
@@
Figure 4: Parameter space in the $a_{\mathrm{in}}$-$a_{\mathrm{out}}$ plane showing which system may undergo resonance capture and advection during the orbital decay of the inner BHB. We consider different types of BHBs as labeled. Assuming the tertiary body is a test mass with a circular orbit, the cross-hatched region indicates dynamically unstable systems, and the gray-shaded region signifies systems with $\gamma\gtrsim 1$, or $\dot{\varpi}_{\mathrm{in}}\gtrsim\dot{\varpi}_{\mathrm{out}}$ (see Eq. 13). The vertical lines refer to the merger time $T_{\mathrm{merger}}=10^{10}$yrs.
The middle and lower panels of Fig. 5 depict four examples of the eccentricity evolution for different $e_{\mathrm{out},0}$ and $\Delta\varpi_{0}$ values, corresponding to different initial phase-space trajectories of the outer binary. For $e_{\mathrm{out},0}=0.1$ (middle panel), we see that $e_{\mathrm{out}}$ experiences resonant excitation and eventually freezes at a finite value; at the same time, $e_{\mathrm{in}}$ experiences ‘extra’ decrease during the resonance advection leading to a decelerated merger event. For $e_{\mathrm{out},0}=0.4$ (lower panel), the eccentricity evolution is sensitive to the initial value of $\Delta\varpi_{0}$ (or the initial phase-space trajectory); the merger undergoes either advection-deceleration or acceleration, as a result of the probabilistic transition during the separatrix crossing.
@@
Summary and discussion.— We have shown that a star orbiting around an inspiraling BHB can experience a distinct resonance capture and advection phenomenon, leading to a significant increase in the eccentricity. This resonance occurs when the apsidal precession rates of the inner and outer orbits are comparable. The resonance advection gives rise to the apsidal alignment, enabling a continuous exchange of angular momentum between the two orbits until the BHB merges. The eccentricity growth requires the inner binary to have an unequal masses and an elliptical orbit, and the outer orbit to have a relatively small eccentricity (see Appendix). For systems with initially large outer eccentricities, the resonance capture occurs probabilistically or may not occur at all. In this study, we have restricted to coplanar configurations. Our preliminary study shows that as long as the mutual inclinations between the inner and outer orbits is small ($\lesssim 30^{\circ}$), similar resonance phenomenon can happen.
The dramatic increase in the outer binary eccentricity $e_{\mathrm{out}}$ has a number of applications. The effective exchange of angular momentum between inner and outer binaries alters the evolution of $e_{\mathrm{in}}$, thereby influencing the hardening rate of the BHBs. Additionally, the notable increase in $e_{\mathrm{out}}$ serves as a distinctive signature for revealing hidden BHBs Liu-Daniel that could be observed by upcoming GW detectors like LISA LISA , TianQin TianQin , Taiji TaiJi and B-DECIGO DECIGO . In some cases, the significant $e_{\mathrm{out}}-$excitation could also result in a sufficiently small pericenter distance for the outer stellar orbit, leading to possibly micro tidal disruption events MTDE .
@@
Appendix A A: Equations of motion
@@
Appendix B B: Resonance capture and advection
@@
Similar to Fig. 2 in the main text, Fig. 7 depicts the dynamic evolution of the vector $\bm{e}_{\mathrm{out}}$ with an initial value of $e_{\mathrm{out},0}=0.6$. As there is no back-reaction from the tertiary on the inner binary, the structure of the phase-space portrait remains the same as Fig. 2, featuring the same contours of constant $\hat{\mathcal{H}}$ and the separatrix for each $\gamma$. However, in contrast to the case with $e_{\mathrm{out},0}\ll 1$, the trajectory of $\bm{e}_{\mathrm{out}}$ here encounters the separatrix around $\gamma\sim 3$ for which zone III already exists. With zone II expanding, both the transition probabilities into zone II and zone III ($P_{\rm I\rightarrow\rm II}$ and $P_{\rm I\rightarrow\rm III}$) are positive (Eqs. 15-16). We find that system initialized with different $\Delta\varpi_{0}$ can undergo either resonance capture or passage, as shown in the upper and lower panels of Fig. 7. The dependence arises because, for a given $e_{\mathrm{out},0}$, the enclosed phase-space areas of $\mathcal{A}_{\rm I(\rm II,III)}$ are different for different $\Delta\varpi_{0}$. Note that for the case of $\Delta\varpi_{0}=0$, the librating evolution trajectory transitions to circulating at $\gamma\simeq 35.1$. This is because zone II starts shrinking in area at this moment, while the phase-space area covered by the evolution trajectory is still conserved. The separatrix crossing becomes unavoidable, and the resonant eccentricity excitation terminates prematurely.
@@
Appendix C C: Apsidal resonance: theoretical analysis

### v2

@@ Abstract
We study the dynamics of a star orbiting a merging black-hole binary (BHB) in a coplanar triple configuration. During the BHB’s orbital decay, the system can be driven across the apsidal precession resonance, where the apsidal precession rate of the stellar orbit matches that of the inner BHB. As a result, the system gets captured into a state of resonance advection until the merger of the BHB, leading to extreme eccentricity growth of the stellar orbit. This resonance advection occurs when the inner binary has a non-zero eccentricity and unequal masses. The resonant driving of the stellar eccentricity can significantly alter the hardening rate of the inner BHB, and produce observational signatures to uncover the presence of nearby merging or merged BHBs.
@@
BHB progenitors, potentially numerous yet undetected, could form an abundant population in stellar multiples (triples or quadruples) Tokovinin ; Raghavan ; Fuhrmann ; Moe-2017 . Observationally, triple systems with massive stars are suggested to be less inclined or nearly coplanar, especially if the tertiary is not so faraway Ransom ; Thompson ; Eisner . If the tertiary companion is a visible object (e.g., a star or a pulsar), its motion can aid in the search for inspiral BHBs Suto-1 ; Suto-2 ; Suto-3 ; Liu-Daniel . In this letter, we focus on the secular dynamics of a star orbiting a merging BHB in a coplanar setup. Beginning with an eccentric BHB and a nearly circular outer stellar orbit, we demonstrate that the system can enter an apsidal resonance and be captured into a ‘resonance advection’ state during the inner BHB’s gravitational radiation driven orbital decay. Once captured, the eccentricity of the star grows continually, attaining a very large value ($\gtrsim 0.9$).
@@
Figure 1: The evolution of a merging BHB with a test-mass tertiary in a coplanar orbital configuration. The masses of BHB are $m_{1}=46.15M_{\odot}$ and $m_{2}=13.85M_{\odot}$. The initial semimajor axes and eccentricities of the inner and outer binaries are $a_{\mathrm{in},0}=0.4\mathrm{AU}$, $e_{\mathrm{in},0}=0.8$, $a_{\mathrm{out}}=5\mathrm{AU}$ and $e_{\mathrm{out},0}=0.1$, and the longitudes of the periapsis of the two orbits are set to be $\varpi_{\mathrm{in},0}=\varpi_{\mathrm{out},0}=0$. The black (inner binary) and orange (outer binary) lines represent the numerical solutions of the DA secular equations derived by the Hamiltonian (Eq. 4). Key moments during the inner BHB evolution are highlighted in the top two panels (gray dots) with the orbital parameter values labeled. The bottom panel displays the apsidal precession rates given by Eqs. (2)-(3).
@@
The top two panels of Fig. 1 show an example of the orbital evolution of a coplanar triple. The system satisfies the dynamical stability criterion Holman . The evolution is obtained by integrating the double-averaged (DA; averaging over both the inner and outer orbital periods) secular equations of motion Liuetal-2015 . All Newtonian effects (up to the octupole order), the leading order PN corrections, and the dissipation due to gravitational wave (GW) are included in the calculations. We see that the BHB undergoes orbital decay and eccentricity circularization due to GW radiation. The merger time of the BHB is approximately given by $T_{\mathrm{merger}}=[5c^{5}a_{\mathrm{in},0}^{4}(1-e_{\mathrm{in},0}^{2})^{7/2}]/[256G^{3}m_{12}^{2}\mu_{\mathrm{in}}]$ Peters-1964 .
@@
and $\dot{\varpi}_{\mathrm{out}}$ is due to the secular perturbation of the inner binary by the outer orbit Liuetal-2015 ; Liu-SMBHB
@@
We see that starting from $\dot{\varpi}_{\mathrm{in}}<\dot{\varpi}_{\mathrm{out}}$ initially, $\dot{\varpi}_{\mathrm{in}}$ increases and $\dot{\varpi}_{\mathrm{out}}$ decreases during orbital decay. The outer eccentricity can be excited once $\dot{\varpi}_{\mathrm{in}}$ catches up to $\dot{\varpi}_{\mathrm{out}}$. This resonance capture is precisely defined in the phase-space diagram (see Fig. 2). Importantly, rather than a single resonance passage, both frequencies evolve in lockstep until the inner BHB merges, resulting in a significant growth of $e_{\mathrm{out}}$. It’s worth noting that the final $e_{\mathrm{out}}$ becomes quite large, and the DA approximation may be challenged. We employ alternative single-averaged equations of motion to assess its validity in the Supplemental material.
Figure 2: Phase-space portraits in terms of the canonical variables ($1-\sqrt{1-e_{\mathrm{out}}^{2}}$ and $\Delta\varpi$) at different stages of the orbital evolution for the system in our fiducial model (refer to Fig. 1). In each panel, $\gamma\equiv\dot{\varpi}_{\mathrm{in}}/\dot{\varpi}_{\mathrm{out}}|_{e_{\mathrm{out}}=0}$ and $\dot{\varpi}_{\mathrm{in}}/\dot{\varpi}_{\mathrm{out}}$ are evaluated by Eqs. (13) and (2)-(3) at the time labeled in Fig. 1. The solid gray lines are the contours of constant $\hat{\mathcal{H}}$. The existence of a separatrix (black line) divides phase-space into circulation zones (zone I and III) and libration zone (zone II). The orange curve represents the numerical solution of the outer binary’s eccentricity vector $\bm{e}_{\mathrm{out}}$, originating at the red dot and ending at the blue dot. To better visualize both evolution trajectory and separatrix, zoom-in subfigures are included in the top-left panels.
@@
By introducing $\Phi_{0}=\Phi|_{a_{\mathrm{in}}=a_{\mathrm{in},0}}$ with $a_{\mathrm{in},0}$ being the initial value of $a_{\mathrm{in}}$, Eq. (8) can be made dimensionless as
@@
Note that apsidal precession resonance and the possibility of exciting eccentricity at resonance haven been well recognized in the planetary dynamics literatures, often in the linear eccentricity regime (e.g., Heppenheimer ; Ward ; Lin ; XuWenrui ; Diego ; Suyubo ). However, the distinct advection effect, sweeping through the full range of canonical momentum, is a unique feature of our present study. Three requirements for the system to end up captured into resonance advection must be satisfied Quillen ; Henrard , including:
@@
iii) Sufficiently small initial eccentricity. Resonance capture requires that the trajectory encounter the separatrix while zone III has not yet appeared. This implies $e_{\mathrm{out},0}\lesssim[15(m_{1}-m_{2})a_{\mathrm{in}}e_{\mathrm{in}}(4+3e_{\mathrm{in}}^{2})]/(8m_{12}a_{\mathrm{out}})$ (see Supplemtary material).
Fig. 3 illustrates the excitation of stellar eccentricity for a range of $(e_{\mathrm{in},0},e_{\mathrm{out},0})$. We find that the cases with $e_{\mathrm{in},0}\gtrsim e_{\mathrm{out},0}$ always give rise to resonant driving of $e_{\mathrm{out}}$, with weak dependence on the initial eccentricities. This behavior can be understood as follows. In the absence of separatrix crossing, as noted before, the phase-space area $\mathcal{A}$ (Eq.14) is an adiabatic invariant. At the time of separatrix crossing, $\mathcal{A}$ is not conserved, and the trajectory can transition into nearby zones Suyubo ; Henrard . For a system initially set with a quasi-circular outer orbit, the trajectory of $\bm{e}_{\mathrm{out}}$ circulates near the bottom part of the phase-space portrait (see the top left panel of Fig. 2). Because there are only two zones at separatrix crossing, the system is guaranteed to be captured into zone II (as in the top panels of Fig. 2). When $0.3\lesssim e_{\mathrm{out}}\lesssim 0.6$, zone III emerges in the transition, and both zone II and III become available. The probabilities of transition to zone $\rm II$ and $\rm III$ are given by Henrard-1982 ; Henrard-1987
@@
where $\mathcal{A}_{\rm I(\rm II,III)}$ is the area of the individual zone. Note that $d\mathcal{A}_{\rm I}+d\mathcal{A}_{\rm II}+d\mathcal{A}_{\rm III}=0$. In our case, both $P_{\rm I\rightarrow\rm II}$ and $P_{\rm I\rightarrow\rm III}$ are positive due to the expansion of zone II and zone III and the contraction of zone I at the separatrix crossing. Whether the system transitions to zone II or zone III depends on the initial $\Delta\varpi_{0}$, which affects the initial phase-space area for a given $e_{\mathrm{out},0}$. Note that when $e_{\mathrm{out},0}=0.6$, the resonant eccentricity excitation terminates prematurely, and $e_{\mathrm{out}}$ reaches around $0.8$. This occurs because after resonance capture, the phase-space area enclosed by the trajectory becomes too large, making separatrix crossing inevitable as zone II begins to shrink (see Supplemental material). If $e_{\mathrm{out},0}$ is even higher, capture into zone II becomes impossible as $P_{\rm I\rightarrow\rm II}$ is negative at the time of separatrix crossing.
Figure 3: The evolution of the outer binary eccentricity as a function of the inner binary semimajor axis. We consider various combinations of ($e_{\mathrm{in},0}$, $e_{\mathrm{out},0}$) with the same parameters as in Fig. 1. Note that for systems with $e_{\mathrm{in},0}=0.8$, the maximum $e_{\mathrm{out},0}$ allowed for a stable triple is $\sim 0.66$ Holman .
@@
Finite $m_{3}$: modified evolution of merging BHBs.— We now consider the ‘BHB$+$star’ system where the tertiary companion has a finite mass $m_{3}=1M_{\odot}$. In this case, the perturbation from the tertiary companion on the inner binary cannot be ignored. Since the outer binary is allowed to ‘gain’ some eccentricity from the inner BHB at resonance, the inner binary may experience a decrease in eccentricity due to the transfer of angular momentum, therefore reducing the orbital decay rate of the BHB. When the initial $e_{\mathrm{out}}$ is large, the inner binary can ‘gain’ some eccentricity, therefore increasing the decay rate.
@@
Figure 4: Parameter space in the $a_{\mathrm{in}}$-$a_{\mathrm{out}}$ plane showing which system may undergo resonance capture and advection during the orbital decay of the inner BHB. We consider different types of BHBs as labeled. Assuming the tertiary body is a test mass with a circular orbit, the cross-hatched region indicates dynamically unstable systems, and the gray-shaded region signifies systems with $\gamma\gtrsim 1$, or $\dot{\varpi}_{\mathrm{in}}\gtrsim\dot{\varpi}_{\mathrm{out}}$. The vertical lines refer to the merger time $T_{\mathrm{merger}}=10^{10}$yrs.
The middle and lower panels of Fig. 5 depict four examples of the eccentricity evolution for different $e_{\mathrm{out},0}$ and $\Delta\varpi_{0}$ values, corresponding to different initial phase-space trajectories of the outer binary. For $e_{\mathrm{out},0}=0.1$ (middle panel), we see that $e_{\mathrm{out}}$ experiences resonant excitation and eventually freezes at a finite value; at the same time, $e_{\mathrm{in}}$ experiences ‘extra’ decrease during the resonance advection, leading to a decelerated merger event. For $e_{\mathrm{out},0}=0.4$ (lower panel), the eccentricity evolution is sensitive to $\Delta\varpi_{0}$ (or the initial phase-space trajectory); the merger undergoes either advection-deceleration or acceleration, as a result of the probabilistic transition during the separatrix crossing.
@@
Summary and discussion.— We have shown that a star orbiting around an inspiraling BHB can experience a distinct resonance capture and advection phenomenon, leading to a significant increase in the eccentricity. This resonance occurs when the apsidal precession rates of the inner and outer orbits are comparable. The resonance advection gives rise to the apsidal alignment, enabling a continuous exchange of angular momentum between the two orbits until the BHB merges. The eccentricity growth requires the inner binary to have an unequal masses and an elliptical orbit, and the outer orbit to have a relatively small eccentricity. For systems with initially large outer eccentricities, the resonance capture occurs probabilistically or may not occur at all. In this study, we have restricted to coplanar configurations. Our preliminary study shows that as long as the mutual inclinations between the inner and outer orbits is small ($\lesssim 30^{\circ}$), similar resonance phenomenon can happen.
The dynamical behavior of the outer binary eccentricity has several implications. The exchange of angular momentum between the inner and outer binaries alters the evolution of $e_{\mathrm{in}}$, thereby influencing the hardening rate of the BHBs. Additionally, the notable increase in $e_{\mathrm{out}}$ serves as a distinctive signature for revealing hidden BHBs Liu-Daniel that could be observed by up-coming GW detectors like LISA LISA , TianQin TianQin , Taiji TaiJi and B-DECIGO DECIGO . In some cases, significant $e_{\mathrm{out}}-$excitation could lead to a small pericenter distance for the outer stellar orbit, resulting significant tidal dissipation or even disruption of the star MTDE .
@@
Supplemental Material
Extreme resonant eccentricity excitation of stars around merging black hole binary
Appendix A Equations of motion
@@
Appendix B Resonance capture and advection
@@
Similar to Fig. 2 in the main text, Fig. 2 depicts the dynamic evolution of the vector $\bm{e}_{\mathrm{out}}$ with an initial value of $e_{\mathrm{out},0}=0.6$. As there is no back-reaction from the tertiary on the inner binary, the structure of the phase-space portrait remains the same as Fig. 2, featuring the same contours of constant $\hat{\mathcal{H}}$ and the separatrix for each $\gamma$. However, in contrast to the case with $e_{\mathrm{out},0}\ll 1$, the trajectory of $\bm{e}_{\mathrm{out}}$ here encounters the separatrix around $\gamma\sim 3$ for which zone III already exists. With zone II expanding, both the transition probabilities into zone II and zone III ($P_{\rm I\rightarrow\rm II}$ and $P_{\rm I\rightarrow\rm III}$) are positive (Eqs. 14-15). We find that system initialized with different $\Delta\varpi_{0}$ can undergo either resonance capture or passage, as shown in the upper and lower panels of Fig. 2. The dependence arises because, for a given $e_{\mathrm{out},0}$, the enclosed phase-space areas of $\mathcal{A}_{\rm I(\rm II,III)}$ are different for different $\Delta\varpi_{0}$. Note that for the case of $\Delta\varpi_{0}=0$, the librating evolution trajectory transitions to circulating at $\gamma\simeq 35.1$. This is because zone II starts shrinking in area at this moment, while the phase-space area covered by the evolution trajectory is still conserved. The separatrix crossing becomes unavoidable, and the resonant eccentricity excitation terminates prematurely.
@@
Appendix C Apsidal resonance: theoretical analysis